# BondMatcher: H-Bond Stability Analysis in Molecular Systems

### v1

@@ Introduction
However, H-bonds are challenging to study as they are weaker than covalent bonds. As a consequence, they can easily rearrange under molecular motions such as vibrations or rotations. Then, the condition of appearance and disappearance of H-bonds under such structural rearrangements is an active research topic, both theoretically and experimentally [28]. Such a flexibility challenges the structural description of H-bonds, which often relies on simple, distance-based, heuristic criteria for establishing their presence, without explicit consideration of the underlying electronic effects [3].
@@ 0.1 Related work
Molecular data analysis and visualization: In recent years, the visualization community has shown a substantial interest in molecular data (as assessed by several surveys [52, 54]), as its complexity presents unique challenges for rendering, exploration and analysis. Several popular software platforms have emerged [47, 84] for the visualization and analysis of molecular systems using ball-and-stick representations. In particular, dedicated rendering techniques have been proposed for the interactive display of these graphs for large-scale molecular systems [15]. Moreover, for dynamic contexts, specialized techniques have also been documented for the tracking of consistent internal frames for these graphs [85, 83]. For even larger systems, molecular surfaces [20, 31, 74] constitute an alternate representation, based on an implicit surface derived from the molecule specification, which facilitates the visualization and analysis of large-scale phenomena (e.g., protein docking [14, 64], cavities [53, 69], pockets [55, 100] and bonding [56]). Specialized techniques for their efficient rendering have also been documented [68, 16, 92]. While they provide a useful approximation of the shape of molecular systems, molecular surfaces only model their geometry, without accounting for the quantum mechanisms governing their structure. For a fine-scale analysis of these mechanisms, several visualization techniques have been established on top of specific scalar descriptors (e.g., the electron density). For instance, approaches based on continuous scatterplots [80] or weighted Voronoi segmentations [59] have been proposed to analyze charge transfers.
@@
Topological Data Analysis: Topological methods [22, 102] have received a considerable attention by the visualization community over the last decades, given their ability to robustly extract structural patterns within scalar data [48]. They introduce a toolbox of data descriptors (persistence diagrams [23, 35], merge [13] and contour trees [12, 33], Reeb graphs [8, 34], or Morse-Smale complexes [77, 81, 42]), where each descriptor captures specific structural relations between the critical points [61] of the input scalar field. The versatility of this toolbox has been demonstrated by the variety of domains where successful applications have been documented, including turbulent combustion [11, 39], material sciences [44, 86], nuclear energy [58], fluid dynamics [51, 65], bioimaging [10, 2], or astrophysics [87, 82]. In particular, in quantum chemistry, these methods enable a robust implementation of QTAIM [7, 91, 67] (with tailored molecular descriptors [32]). At a conceptual level, our work deals with the study of the variability of a specific topological descriptor (the extremum graph, Sec. 3), within an ensemble of scalar fields, which is a topic which also received a significant attention in recent years. For instance, Athawale et al. [4] introduce a framework for the pointwise analysis of the geometrical variability of separatrices in the 2D Morse complex, based on an overlap-driven matching of mandatory critical points [37]. Also, general-purpose metrics [63] and tracking algorithms [21] have been introduced for extremum graphs. In contrast to the above methods, our work focuses on studying the global structural variability of separatrices in a specific application context (ensembles of 3D electron density fields), which comes with domain specific hypotheses, which our work specifically exploits, resulting in a simple and effective algorithm.
@@ 1.1 Hydrogen bonds
Figure 1: Main concepts from the Quantum Theory of Atoms in Molecule (QTAIM) on a 2D slice of the electron density for a water molecule. Left: Maxima (red spheres) of the electron density (gray gradient) represent atoms, while saddles (white spheres) represent bond critical points. Right: Bond critical points are connected to atoms via bond paths (black cylinders). In QTAIM, two atoms that are considered chemically bonded are linked by a bond path [60]. Each atom is associated with its basin (colored regions, in pink, light gray and dark gray), whose boundary forms an interatomic surface (red curves on this 2D sliced view).
@@ 1.3 Topological Data Analysis
Input data: The input data is typically provided by the DFT computation as a scalar field $\rho:\mathcal{C}\rightarrow\mathbb{R}$, valued on the vertices of a cubical complex $\mathcal{C}$ (i.e., a three-dimensional regular grid). In practice, for implementation genericity purposes, $\mathcal{C}$ is often triangulated (via an on-the-fly emulation [91]) into a simplicial complex $\mathcal{K}$ (typically, with the Freudenthal triangulation [45, 50]). This yields a piecewise linear (PL) scalar field $\rho:\mathcal{K}\rightarrow\mathbb{R}$, valued on the vertices of $\mathcal{K}$ and linearly interpolated on the simplices of higher dimensions. At this stage, the input $\rho$ is assumed to be injective on the vertices of $\mathcal{K}$, which is enforced in practice via symbolic perturbation [24].
@@
Discrete gradient field: Given the above input formalization, one can leverage discrete Morse theory (DMT) [27] to design robust, combinatorial algorithm for Topological Data Analysis (as illustrated in Fig. 2). For this, the notion of discrete gradient field can be introduced as follows. Let $\{\sigma_{i}<\sigma_{i+1}\}$ be a discrete vector. It is formed by a pair of simplices of $\mathcal{K}$, such that $\sigma_{i}$ is an $i$-simplex and $\sigma_{i+1}$ is a cofacet of $\sigma_{i}$ (see the vertex-edge vectors in black and the edge-triangle vectors in red in Fig. 2, top). $\sigma_{i+1}$ is usually referred to as the head of the vector, while $\sigma_{i}$ is its tail. Then, a discrete vector field on $\mathcal{K}$ is defined as a collection of pairs $\{\sigma_{i}<\sigma_{i+1}\}$, such that each simplex of $\mathcal{K}$ is included in at most one discrete vector. A simplex involved in no discrete vector is called a critical simplex (larger spheres in Fig. 2). The notion of v-path introduces a discrete analog to the concept of integral line. It is defined as a sequence of $k$ discrete vectors $\big{\{}\{\sigma_{i}^{1}<\sigma_{i+1}^{1}\},\dots,\big{\{}\{\sigma_{i}^{k}<\sigma_{i+1}^{k}\}\big{\}}$ such that (i) $\sigma_{i}^{j}\neq\sigma_{i}^{j+1}$ (i.e., the tails of two consecutive vectors are distinct) and (ii) $\sigma_{i}^{j+1}<\sigma_{i+1}^{j}$ (i.e., the tail of a vector in the sequence is a face of the head of the previous vector in the sequence) for any $1<j<k$. A v-path is said to terminate at a critical simplex $\sigma_{i}$ if $\sigma_{i}$ is a facet of the head of its last vector. By symmetry, a v-path is said to start at a critical simplex $\sigma_{i+1}$ if $\sigma_{i+1}$ is a cofacet of the tail of its first vector. Then, we call the stable set of a critical simplex $\sigma_{i}$ the collection of v-paths which terminate in $\sigma_{i}$ (Fig. 2). By symmetry, we call the unstable set of $\sigma_{i+1}$ the collection of v-paths which start in $\sigma_{i+1}$ (cylinders in Fig. 2, bottom). Finally, a discrete vector field for which all v-paths are loop-free is called a discrete gradient field.
@@
Several algorithms have been documented for the computation of such a discrete gradient field from an input scalar field [43, 77, 81]. Then, the critical simplices are discrete analogs to the notion of critical points from the smooth setting [61], with their dimension $i$ matching with the smooth notion of critical index.
@@
Morse complex: Given a discrete gradient field, the Morse complex is the complex formed by the collection of stable sets (Fig. 2, bottom). In practice, to cope with noise or numerical inaccuracies, the Morse complex is often simplified, either in a post-process [38, 40] or in a pre-process [36], by modifying the discrete gradient field prior to computing the Morse complex. In particular, a pair of critical simplices $\sigma_{i}$ and $\sigma_{i+1}$ can be simplified together if there exists a unique v-path connecting them, by simply reversing the discrete vectors along the v-path [27]. Then, the Morse complex can be progressively simplified, by iteratively reverting v-paths connecting pairs of critical simplices (Fig. 2, from left to right). In practice, several criteria can be considered for selecting a v-path for reversal. For instance, topological persistence [23, 22] is an established, general-purpose importance measure, that is well studied in Topological Data Analysis. Overall, this combinatorial approach enables a robust and straightforward implementation of QTAIM [7, 91, 67, 9]. Specifically, since in DMT discrete vectors point down [27], the QTAIM basins (respectively the interatomic surfaces) correspond to the stable sets of the minima (respectively $1$-saddles) of the opposite of the electron density. Moreover, the bond paths are given by the unstable sets of $1$-saddles of this opposite density.
@@ 2 Electron Density Ensemble Generation
This section describes the generation of the ensemble datasets considered in our case studies (Sec. 5.3 and Sec. 5.4).
Figure 3: Extremum graphs of electron density, at various stages of bond graph extraction, for a Prism water hexamer (transparent: isosurface for $\rho^{\prime}=-0.1$). (a) The presence of spurious minima (red sphere) requires topological simplification [57]. Minima are colored by topological persistence (Sec. 1.3). In this example, the spurious minimum (red) has a persistence of $4\times 10^{-4}$, while the first non-spurious one has a persistence of $3.6\times 10^{-2}$ (two order of magnitude gap). (b) The presence of spurious pairs of $1$ and $2$-saddles (orange and black spheres respectively) requires a step of simplification, achieved by v-path reversals, ordered by function value difference. In this example, the last spurious saddle-saddle v-path exhibits a function difference of $4\times 10^{-5}$, while the first non-spurious one has a difference of $5\times 10^{-3}$ (two order of magnitude gap). (c) The simplified extremum graph is represented with its nodes and arcs colored by $\rho^{\prime}$ (nodes: value at the minimum, arcs: value at the saddle), enabling a clear distinction between oxygen (red) and hydrogen (pink) atoms as well as between covalent (pink) and non-covalent (gray to black) bonds. (d) The resulting bond graph represents the identified atoms (red: oxygen, white: hydrogen) and bond types (dark gray: covalent, light gray: H-bond).
@@ 2.1 Water cluster selection
The study of water clusters, (H2O)${}_{n_{m}}$, is essential for the understanding of the structure of liquid water and ice [98, 29]. It provides key insights into the mechanisms of many processes, such as water aggregation, microsolvation [26, 78, 46, 99] or proton transfer [96].
For a given number of molecules $n_{m}$, water clusters can exhibit distinct isomeric forms (i.e., geometrical arrangements), as each water molecule can partcipate in four H-bonds. Each isomer has a specific hydrogen bond network (i.e., set of H-bonds), resulting in unique chemical properties. Moreover, isomers with comparable energies can easily rearrange into one another through a process called interconversion [28].
Water hexamers ($n_{m}=6$) are prominent examples of water clusters [29], as they constitute the smallest three-dimensional cluster whose non-quasiplanar isomers (the Book, Cage and Prism) exhibit lower energies than quasiplanar isomers (such as the Ring). For this reason, (H2O)6 is often referred to in the literature as “the smallest drop of water” [94, 66]. High resolution spectroscopy data as well as state-of-the-art quantum calculations [79, 71, 94] indicate that the Prism, the Cage and the Book are the isomers of (H2O)6 with the lowest energies. Therefore, they likely coexist, even at low temperatures. Thus, we focus our study on these three isomers as well as the Ring isomer, which is the quasiplanar structure with the lowest energy.
@@ 2.2 Tunneling pathways
To evaluate the chemical relevance of our approach, we first study the Prism isomer under the effect of proton tunneling - a phenomenon in which a molecule in its equilibrium state rearranges to an equivalent (i.e., isoenergetic) structure, for example by permuting equivalent atoms [72]. Proton tunneling in the Prism isomer of (H2O)6 was observed experimentally and confirmed by calculations [76, 93, 18]. Specifically, far-infrared vibration-rotation tunneling (FIR-VRT) spectroscopy observations report characteristic splitting patterns. These can be explained by two distinct tunneling mechanisms, associated to specific rotations of water molecules: (i) An anti-geared rotation, breaking one H-bond, and (ii) a geared motion, simultaneously breaking two H-bonds. Each of these two motions can be modeled by a specific pathway (i.e., a temporal sequence of transitory structures between the two equivalent states), which can be estimated with the ring-polymer molecular dynamics (RPMD) [75]. For each step $\alpha_{k}\in(0,255)$ of the pathways reported in [76], the electron density was computed with a non-relativistic DFT method [49], with the PBE0 functional [1] and the TZ2P basis, as implemented in the ADF software [89] and facilitated through PyADF[25]. The computations were carried at on a $256^{3}$ regular grid, with a spacing of $0.05$ Angstrom (Å) between grid points.

### v2

@@ Introduction
Nevertheless, H-bonds are challenging to study as they are weaker than covalent bonds. As a consequence, they can easily rearrange under molecular motions such as vibrations or rotations. Then, the condition of appearance and disappearance of H-bonds under such structural rearrangements is an active research topic, both theoretically and experimentally [24]. Such a flexibility challenges the structural description of H-bonds, which often relies on simple, distance-based, heuristic criteria for establishing their presence, without explicit consideration of the underlying electronic effects [1].
@@ 0.1 Related work
Molecular data analysis and visualization: In recent years, the visualization community has shown a substantial interest in molecular data (as assessed by several surveys [44]), as its complexity presents unique challenges for rendering, exploration and analysis. Several popular software platforms have emerged [40, 73] for the visualization and analysis of molecular systems using ball-and-stick representations. In particular, dedicated rendering techniques have been proposed for the interactive display of these graphs for large-scale molecular systems [12]. Moreover, for dynamic contexts, specialized techniques have also been documented for the tracking of consistent internal frames for these graphs [74, 72]. For even larger systems, molecular surfaces [16] constitute an alternate representation, based on an implicit surface derived from the molecule specification, which facilitates the visualization and analysis of large-scale phenomena (e.g., protein docking [11, 54], cavities [45, 59], pockets [46, 88] and bonding [47]). Specialized techniques for their efficient rendering have also been documented [58, 80]. While they provide a useful approximation of the shape of molecular systems, molecular surfaces only model their geometry, without accounting for the quantum mechanisms governing their structure. For a fine-scale analysis of these mechanisms, several techniques have been established on top of specific scalar descriptors (e.g., the electron density). For instance, approaches based on continuous scatterplots [69] or Voronoi segmentations [49] have been proposed for charge transfer analysis.
@@
Topological Data Analysis: Topological methods [19, 90] have received a considerable attention by the visualization community over the last decades, given their ability to robustly extract structural patterns within scalar data [41]. They introduce a toolbox of data descriptors (persistence diagrams [20, 29], merge and contour trees [10, 27], Reeb graphs [6, 28], or Morse-Smale complexes [66, 70, 35]), where each descriptor captures specific structural relations between the critical points [51] of the input scalar field. The versatility of this toolbox has been demonstrated by the variety of domains where successful applications have been documented, including turbulent combustion [9, 32], material sciences [36, 75], fluid dynamics [43, 55], bioimaging [8], or astrophysics [71]. In particular, in quantum chemistry, these methods enable a robust implementation of QTAIM [5, 79, 57] (with tailored molecular descriptors [37]). At a conceptual level, our work deals with the study of the variability of a specific topological descriptor (the extremum graph [17], Sec. 3), within an ensemble of scalar fields, which is a topic which also received a significant attention in recent years. For instance, Athawale et al. [2] introduce a framework for the pointwise analysis of the geometrical variability of separatrices in the 2D Morse complex, based on an overlap-driven matching of mandatory critical points [31]. Also, general-purpose metrics [53] and tracking algorithms [18] have been introduced for extremum graphs. In contrast to the above methods, our work focuses on studying the global structural variability of separatrices in a specific application context (ensembles of 3D electron density fields), which comes with domain specific hypotheses, which our work specifically exploits, resulting in a simple and effective algorithm.
@@ 1.1 Hydrogen bonds
Figure 1: Main concepts from the Quantum Theory of Atoms in Molecule (QTAIM) on a 2D slice of the electron density for a water molecule. Left: Maxima (red spheres) of the electron density (gray gradient) represent atoms, while saddles (white spheres) represent bond critical points. Arrows illustrate extreme vibrational displacement vectors (mode 00, see Sec. 2.3). Right: Bond critical points are connected to atoms via bond paths (black cylinders). In QTAIM, two atoms that are considered chemically bonded are linked by a bond path [50]. Each atom is associated with its basin (colored regions, pink, light gray and dark gray), whose boundary forms an interatomic surface (red curves on this 2D slice).
@@ 1.3 Topological Data Analysis
Input data: The input data is typically provided by the DFT computation as a scalar field $\rho:\mathcal{C}\rightarrow\mathbb{R}$, valued on the vertices of a cubical complex $\mathcal{C}$ (i.e., a three-dimensional regular grid). In practice, for implementation genericity purposes, $\mathcal{C}$ is often triangulated (via an on-the-fly emulation [79]) into a simplicial complex $\mathcal{K}$ (typically, with the Freudenthal triangulation [38]). This yields a piecewise linear (PL) scalar field $\rho:\mathcal{K}\rightarrow\mathbb{R}$, valued on the vertices of $\mathcal{K}$ and linearly interpolated on the simplices of higher dimensions. At this stage, the input $\rho$ is assumed to be injective on the vertices of $\mathcal{K}$, which is enforced in practice via symbolic perturbation [21].
@@
Discrete gradient field: Given the above input formalization, one can leverage discrete Morse theory (DMT) [23] to design robust, combinatorial algorithms for Topological Data Analysis (as illustrated in Fig. 2). For this, the notion of a discrete gradient field can be introduced as follows. Let $\{\sigma_{i}<\sigma_{i+1}\}$ be a discrete vector. It is formed by a pair of simplices of $\mathcal{K}$, such that $\sigma_{i}$ is an $i$-simplex and $\sigma_{i+1}$ is a cofacet of $\sigma_{i}$ (see the vertex-edge vectors in black and the edge-triangle vectors in red in Fig. 2, top). $\sigma_{i+1}$ is usually referred to as the head of the vector, while $\sigma_{i}$ is its tail. A discrete vector field on $\mathcal{K}$ is defined as a collection of pairs $\{\sigma_{i}<\sigma_{i+1}\}$, such that each simplex of $\mathcal{K}$ is included in at most one discrete vector. A simplex involved in no discrete vector is called a critical simplex (larger spheres in Fig. 2). The notion of a v-path introduces a discrete analog to the concept of integral line. It is defined as a sequence of $k$ discrete vectors $\big{\{}\{\sigma_{i}^{1}<\sigma_{i+1}^{1}\},\dots,\big{\{}\{\sigma_{i}^{k}<\sigma_{i+1}^{k}\}\big{\}}$ such that (i) $\sigma_{i}^{j}\neq\sigma_{i}^{j+1}$ (i.e., the tails of two consecutive vectors are distinct) and (ii) $\sigma_{i}^{j+1}<\sigma_{i+1}^{j}$ (i.e., the tail of a vector in the sequence is a face of the head of the previous vector in the sequence) for any $1<j<k$. A v-path is said to terminate at a critical simplex $\sigma_{i}$ if $\sigma_{i}$ is a facet of the head of its last vector. By symmetry, a v-path is said to start at a critical simplex $\sigma_{i+1}$ if $\sigma_{i+1}$ is a cofacet of the tail of its first vector. We call the stable set of a critical simplex $\sigma_{i}$ the collection of v-paths which terminate in $\sigma_{i}$ (Fig. 2). By symmetry, we call the unstable set of $\sigma_{i+1}$ the collection of v-paths which start in $\sigma_{i+1}$ (cylinders in Fig. 2, bottom). Finally, a discrete vector field for which all v-paths are loop-free is called a discrete gradient field.
@@
Several algorithms have been documented for the computation of such a discrete gradient field from an input scalar field [66, 70]. In that context, the critical simplices are discrete analogs to the notion of critical points from the smooth setting [51], with their dimension $i$ matching with the smooth notion of critical indices.
@@
Morse complex: Given a discrete gradient field, the Morse complex is the complex formed by the collection of stable sets (Fig. 2, bottom). In practice, to cope with noise or numerical inaccuracies, the Morse complex is often simplified, either in a post-process [33] or in a pre-process [30], by modifying the discrete gradient field prior to computing the Morse complex. In particular, a pair of critical simplices $\sigma_{i}$ and $\sigma_{i+1}$ can be simplified together if there exists a unique v-path connecting them, by simply reversing the discrete vectors along the v-path [23]. Globally, the Morse complex can be progressively simplified, by iteratively reverting v-paths connecting pairs of critical simplices (Fig. 2, from left to right). In practice, several criteria can be considered for selecting a v-path for reversal. For instance, topological persistence [20, 19] is an established, general-purpose importance measure, that is well studied in Topological Data Analysis. Overall, this combinatorial approach enables a robust and straightforward implementation of QTAIM [5, 79, 57, 7]. Specifically, since in DMT discrete vectors point down [23], the QTAIM basins correspond to the stable sets of the minima of the opposite of the electron density, and the interatomic surfaces correspond to the stable sets of its $1$-saddles. Also, the bond paths are given by the unstable sets of $1$-saddles of this opposite density.
@@ 2 Electron Density Ensemble Generation
This section describes the generation of the ensemble datasets considered in our case studies (Sec. 5.2 and Sec. 5.3). These ensembles model, for selected water clusters (Sec. 2.1), proton tunneling (Sec. 2.2) and molecular vibrations (Sec. 2.3). These are two important phenomena naturally occurring in chemical processes, which can trigger isomerization – transforming a molecule into another, with identical composition, but with a distinct spatial arrangement of atoms, resulting in distinct chemical properties. Isomerization plays a crucial role in many areas of chemistry (e.g., drug design) and understanding the conditions favoring its occurrence is an important challenge in quantum chemistry.
Figure 3: Extremum graphs of opposite electron density ($\rho^{\prime}\in(-126,0)$), at various stages of bond graph extraction, for a Prism water hexamer (transparent: isosurface for $\rho^{\prime}=-0.1$). (a) The presence of spurious minima (red sphere) requires topological simplification [48]. Minima are colored by topological persistence (Sec. 1.3). In this example, the spurious minimum (red) has a persistence of $4\times 10^{-4}$, while the first non-spurious one has a persistence of $3.6\times 10^{-2}$ (two order of magnitude gap). (b) The presence of spurious pairs of $1$ and $2$-saddles (orange and black spheres respectively) requires a step of simplification, achieved by v-path reversals, ordered by function value difference. Here, the last spurious saddle-saddle v-path exhibits a function difference of $4\times 10^{-5}$, while the first non-spurious one has a difference of $5\times 10^{-3}$ (two order of magnitude gap). (c) The simplified extremum graph is represented with its nodes and arcs colored by $\rho^{\prime}$ (nodes: value at the minimum, arcs: value at the saddle), enabling a clear distinction between oxygen (red) and hydrogen (pink) atoms as well as between covalent (pink) and non-covalent (white to black) bonds. (d) The resulting bond graph represents the identified atoms (red: oxygen, white: hydrogen) and bond types (dark gray: covalent, light gray: H-bond).
@@ 2.1 Water cluster selection
The study of water clusters, (H2O)${}_{n_{m}}$, is essential for the understanding of the structure of liquid water and ice [85, 25]. It provides key insights into the mechanisms of many processes, such as water aggregation, microsolvation [67, 39, 87] or proton transfer [83].
For a given number of molecules $n_{m}$, water clusters can exhibit distinct isomeric forms (i.e., geometrical arrangements), as each water molecule can partcipate in four H-bonds. Each isomer has a specific hydrogen bond network (i.e., set of H-bonds), resulting in unique chemical properties. Moreover, isomers with comparable energies can easily rearrange into one another through a process called interconversion [24]. In our work, we focus on these relatively simple yet chemically important systems due to their extensive documentation and ubiquity in chemistry. For example, several studies document proton tunneling (Sec. 2.2) or molecular vibration (Sec. 2.3) for these clusters, hence helping validating our approach. Also, their geometric scale, between micro- and macro-systems, typically makes them good candidates in chemistry for investigating novel models or experimental protocoles, before considering more complex systems (see Sec. 6).
Specifically, we focus on water hexamers ($n_{m}=6$), which are prominent examples of water clusters [25], as they constitute the smallest three-dimensional cluster whose non-quasiplanar isomers (the Book, Cage and Prism) exhibit lower energies than quasiplanar isomers (such as the Ring). For this reason, (H2O)6 is often referred to in the literature as “the smallest drop of water” [56]. High resolution spectroscopy as well as state-of-the-art quantum calculations [68, 61] indicate that the Prism, the Cage and the Book are the isomers of (H2O)6 with the lowest energies. Consequently, they likely coexist, even at low temperatures. Therefore, we focus our study on these three isomers as well as the Ring isomer, which is the quasiplanar structure with the lowest energy.
@@ 2.2 Tunneling pathways
To evaluate the chemical relevance of our approach, we first study the Prism isomer under the effect of proton tunneling - a phenomenon in which a molecule in its equilibrium state rearranges to an equivalent (i.e., isoenergetic) structure, for example by permuting equivalent atoms [62]. Proton tunneling in the Prism isomer of (H2O)6 was observed experimentally and confirmed by calculations [65, 81, 14]. Specifically, far-infrared vibration-rotation tunneling (FIR-VRT) spectroscopy observations report characteristic splitting patterns. These can be explained by two distinct tunneling mechanisms, associated to specific rotations of water molecules: (i) An anti-geared rotation, breaking one H-bond, and (ii) a geared motion, simultaneously breaking two H-bonds. Each of these two motions can be modeled by a specific pathway (i.e., a temporal sequence of transitory structures between the two equivalent states), which can be estimated with the ring-polymer molecular dynamics (RPMD) approach [64]. For each step $\alpha_{k}\in(0,255)$ of the pathways reported in [65], the electron density was computed with a non-relativistic DFT method [42], with the PBE0 functional [86] and the TZ2P basis, as implemented in the ADF software [77] and facilitated through PyADF[22]. The computations were carried out on a $256^{3}$ regular grid, with a sample spacing of $0.05$ Angstrom (Å).